# ASER: Activation Smoothing and Error Reconstruction for Large Language Model Quantization

### v1

@@ Introduction
Quantization involves converting the model’s weights and activations from high-precision floating-point numbers to lower-precision integers, whose primary goal is alleviating the computational and memory requirements without significantly compromising the model’s performance, enabling edge deployment. Take post training quantization (PTQ) as an example, current advancements try to smooth (Xiao et al. 2023), rotate (Liu et al. 2024) or transform tensor representations in models, showing promise of achieving quantization-friendly data range and remarkable compression rate. However, when it comes to low-bit quantization, even with the application of the aforementioned techniques, the error introduced by the quantized model becomes evident, leading to notable degradation in the evaluation metrics.
As for typical quantization, the optimization objective can be formulated as minimizing $\|\textbf{WX}-\textbf{W}_{q}\textbf{X}\|_{F}$, where W and $_{q}=Q(\textbf{W})$ W ) are original model weight and its quantized one respectively, and X is the activation. We empirically find this inevitable quantization error has low-rank property, which prompts us to consider using a method similar to LoRA (Hu et al. 2021) to reconstruct this error. Specifically, the activation-weight quantization loss exhibits distinct low-rank property, with its singular value distribution featuring a small number of high values and a long-tail bulk of low ones, which varies both intra and inter Transformer blocks. Furthermore, we find the channels that produce the major error are consistent with the existence of the outliers.
@@
We propose ASER (Activation Smoothing and Error Reconstruction), a low-rank compensation algorithm designed to enhance the efficacy of quantized models. Our method dynamically assesses the low-rank properties during the quantization process, whitens the relationship between quantization loss and singular values based on Cholesky factorization. Then, ASER identifies and smoothes these error-inducing outliers for separate compensation. Finally, we use LoRA-style skinny matrices to compensate the quantization error, with little computation overhead. The overall framework is shown in Figure 1. This lightweight yet powerful mechanism ensures robust performance of our quantized models, and this approach is orthogonal to any particular weight quantization method. Our experiments demonstrate that ASER remarkably recovers the quantization error, where W4A8 per-channel quantized model performs nearly equal capability related to the half-precision reference model. This work makes the following contributions:
  • We formulate the optimization objective of minimizing the discrepancy in model outputs, analyze the characteristics of quantization errors in LLMs quantization.
@@
$L_{i}$ $=\|(\textbf{E}_{q}-\tilde{\textbf{E}}^{i}_{q})\textbf{X}\|_{F}$ (8)
$=\|\sigma_{i}\textbf{u}_{i}\textbf{v}_{i}^{\top}\textbf{S}^{-1}\textbf{X}\|_{F}$
@@
=$\overset{(ii)}{=}\sigma_{i}\operatorname{Tr}(\textbf{u}_{i}\textbf{v}_{i}^{\top}\textbf{v}_{i}\textbf{u}_{i}^{\top})^{\frac{1}{2}}$
=$\overset{(iii)}{=}\sigma_{i},$
@@
Note that the error reconstruction technique is orthogonal to the weight-only quantization, which means ASER can be conducted based on the existing weight-only quantization methods, not limited to RTN. The pseudo-code of our algorithm is as follows:
@@
13:     Compute S by Cholesky decomposition on $\textbf{M}^{-1}\textbf{X}$: $(\textbf{S}^{-1}\textbf{M}^{-1}\textbf{X})(\textbf{S}^{-1}\textbf{M}^{-1}\textbf{X})^{\top}=\textbf{I}$
14:     Compute SVD: $\textbf{U}\mathbf{\Sigma}\textbf{V}^{\top}=\textbf{E}_{q}^{l}\textbf{S}$
@@
Experimental Details To ensure a fair comparison in our assessment, we standardize the usage of the calibration dataset for all quantization methods. All quantization is conducted by per-channel for weights and per-token for activations. We consistently employ a sample size of $128$ and maintain a uniform token sequence length of $2048$ for calibration throughout the evaluation process. All experiments are conducted on 4×4$\times$4 × NVIDIA A100 GPU.
@@ Conclusion
In this work, from the perspective of model compression, we formulate and analyze the quantization error. Our study reveals that the error exhibits low-rank property, whose major components originate from the combined effect of outliers in both activations and weights. In response to these findings, we propose ASER, which utilizes whitening SVD to compensate for the integral error (Error Reconstruction), followed by analyzing and extracting outliers (Activation Smoothing). Experimental results demonstrate that ASER outperforms existing quantization algorithms on typical large language models in various quantization setups, with negligible theoretical overhead.
Appendix A Appendix
Numerical Effect of Activation Smoothing
The idea of activation smoothing is to extract the outliers composed by both weights and activations, and transfer the difficulty of activation quantization, then the outliers part will be construct within the quantization error reconstruction. In Fig. 7, we demonstrate that, in the numerical perspective, the activation range is significantly smoothed, while the outliers transfered into the weight are extracted by $W_{o}$.
Figure 7: The effect of activation smoothing. The figures on the left and right respectively represent $A$ and $W$ before and after smoothing, where $WM=W_{s}+W_{o}$. The data comes from the first layer in Qwen1.5-7B.
Effect of Rank Selection Strategy
As mentioned before, we observe the different effective dimentionality of quantization error across the layers in LLM. When it comes to the construction of LoRA-style matrix to compensate for the error, we design the strategy to dynamically select the rank configuration. As Fig. 8 shows, when using $\alpha$ ranging from $0.015$ to $0.1$, we employ different ranks across layers, whose trends are relatively consistent within the same layer across the Transformer blocks. The distribution also conforms to the effective rank in our analysis of quantization error.
Figure 8: The effect of rank selection. The rank is set by different threshold $\alpha$ in different layers from LLaMA3-8B.
More Evaluation Results
We extensively evaluate ASER on models of various scales and types, and on different datasets, to demonstrate its generalization ability. Specifically, we supplements the weight-only quantization on LLaMA3-8B (Table 5). Compared to the state-of-the-art approaches GPTQ and AWQ, ASER achieves competitive results both in causal language modeling perplexity and commonsense tasks accuracy. Similar results of weight-only quantization setup are shown in LLaMA2-13B (Table 6).
Note that, with activation smoothing, the advantage of ASER lies in its quantization of activations. Therefore, we emphasize testing on act-and-weight quantization. Besides LLaMA2-13B (Table 6), larger models are involved in the evaluation. As for Qwen-14B (Table 7), Qwen1.5-32B (Table 8), ASER shows close average accuracy related to the fp16 baseline model in the W4A8 per-channel quantization setup.
WikiText2 C4 PTB ARC-e ARC-c MMLU Hella PIQA Avg.
LLaMA3-8B 16 16 6.14 10.75 9.91 89.77 79.66 37.85 48.47 68.55 64.86
RTN 4 16 10.21 17.86 15.09 79.01 60.68 12.89 30.85 55.82 47.85
GPTQ 4 16 7.91 13.81 11.86 79.89 66.78 26.43 36.07 61.32 54.10
AWQ 4 16 7.86 13.70 11.74 85.36 72.88 20.47 36.26 60.99 55.19
ASER (w/o AS) 4 16 7.34 12.74 11.20 86.77 74.24 28.54 31.51 58.87 55.99
ASER (w/ AS) 4 16 7.44 12.94 11.40 86.95 72.88 31.20 29.39 60.55 56.19
Table 6: Evaluation results of post-training quantization on LLaMA2-13B. Best results are marked in bold.
WikiText2 C4 PTB ARC-e ARC-c MMLU Hella PIQA Avg.
LLaMA2-13B 16 16 4.88 7.55 34.42 68.08 48.47 45.42 31.92 55.88 49.95
RTN 4 16 5.39 8.21 46.13 64.20 46.78 37.21 29.36 57.62 47.03
GPTQ 4 16 5.22 8.05 74.85 63.32 52.88 42.40 31.02 53.81 48.69
AWQ 4 16 5.33 8.14 50.17 59.96 42.37 40.27 28.44 55.71 45.35
ASER (w/o A.S.) 4 16 5.15 7.85 39.09 68.43 48.81 47.30 31.96 54.90 50.28
ASER (w/ A.S.) 4 16 5.12 7.85 38.65 67.37 50.51 44.82 31.02 55.88 49.92
LLM.int4() 4 8 5.46 8.30 48.77 63.49 44.75 36.39 28.75 57.45 46.17
SmoothQuant 4 8 6.10 9.26 52.60 39.66 22.75 30.03 25.26 50.44 33.63
LoRC 4 8 5.28 8.03 38.45 65.08 50.17 44.83 30.11 57.07 49.45
L2QER 4 8 5.21 7.95 39.57 67.20 48.81 45.19 30.09 56.64 49.59
ASER (w/o A.S.) 4 8 5.21 7.92 40.45 68.08 48.14 47.81 31.22 58.16 50.68
ASER (w/ A.S.) 4 8 5.16 7.89 39.44 67.02 49.83 45.08 29.83 57.56 49.86
Table 7: Evaluation results of post-training quantization on Qwen-14B. Best results are marked in bold.
Method #W #A Accuracy()($\uparrow$)( ↑ )
ARC-e ARC-c Hella PIQA Avg.
Qwen-14B 16 16 92.24 83.05 80.42 73.88 82.40
LLM.int4() 4 8 86.77 79.32 41.31 52.99 65.10
SmoothQuant 4 8 55.73 45.08 31.33 49.18 45.33
LoRC 4 8 92.06 84.41 65.31 67.08 77.22
L2QER 4 8 91.01 83.05 71.68 69.42 78.79
ASER (w/o A.S.) 4 8 91.36 82.03 71.94 70.08 78.85
ASER (w/ A.S.) 4 8 91.71 82.03 74.15 70.84 79.68
Table 8: Evaluation results of post-training quantization on Qwen1.5-32B. Best results are marked in bold.
Method #W #A Accuracy()($\uparrow$)( ↑ )
ARC-e ARC-c Hella PIQA Avg.
Qwen1.5-32B 16 16 84.30 71.53 16.26 56.96 57.26
LLM.int4() 4 8 68.61 51.19 6.88 44.78 42.86
SmoothQuant 4 8 69.14 53.90 6.10 47.17 44.08
LoRC 4 8 81.83 69.49 16.21 51.69 54.81
L2QER 4 8 82.89 70.17 15.81 54.35 55.81
ASER (w/o A.S.) 4 8 85.19 69.15 14.66 54.84 55.96
ASER (w/ A.S.) 4 8 85.01 73.56 15.08 54.52 57.04

### v2

@@ Introduction
Quantization involves converting the model’s weights and activations from floating-point numbers to lower-precision integers, whose primary goal is alleviating the computational and memory requirements without significantly compromising the model’s performance, enabling edge deployment. Take post training quantization (PTQ) as an example, current methods try to smooth (Xiao et al. 2023), rotate (Liu et al. 2024) or transform representations in models, showing promise of achieving quantization-friendly data range and remarkable compression rate. However, when it comes to low-bit quantization, even with the application of the aforementioned techniques, the error introduced by the quantized model becomes evident, resulting in notable degradation.
For typical quantization, the optimization objective can be formulated as minimizing $\|\textbf{WX}-\textbf{W}_{q}\textbf{X}\|_{F}$, where W and $_{q}=Q(\textbf{W})$ W ) are original model weight and its quantized one respectively. We empirically find this inevitable quantization error has low-rank property, whose singular value distribution featuring a small number of high values and a long-tail bulk of low ones. This varies both intra and inter Transformer blocks, which prompts us to consider using a module similar to LoRA (Hu et al. 2021) to reconstruct this error. Furthermore, we find the channels produce the major error are consistent with the existence of the outliers.
@@
We propose ASER (Activation Smoothing and Error Reconstruction), a low-rank compensation algorithm designed to enhance the efficacy of quantized models. Our method assesses the low-rank properties by whitening the relationship between quantization loss and its singular values based on Cholesky factorization. Then, ASER identifies these error-inducing outliers for smoothing. Finally, LoRA-style matrices are used to reconstruct the quantization error, with little computation overhead. The overall framework is shown in Figure 1. This lightweight yet powerful mechanism ensures robust performance of quantized models, which is orthogonal to any particular weight quantization method. Our experiments demonstrate ASER remarkably recovers the quantization error, where W4A8 per-channel quantized model performs nearly equal capability compared to the half-precision reference. This work makes the following contributions:
  • We formulate the optimization objective of minimizing the discrepancy in model outputs, and analyze the characteristics of quantization errors in LLMs quantization.
@@
$L_{i}$ $=\|(\textbf{E}_{q}-\tilde{\textbf{E}}^{i}_{q})\textbf{X}\|_{F}=\|\sigma_{i}\textbf{u}_{i}\textbf{v}_{i}^{\top}\textbf{S}^{-1}\textbf{X}\|_{F}$ (8)
@@
=$\overset{(ii)}{=}\sigma_{i}\operatorname{Tr}(\textbf{u}_{i}\textbf{v}_{i}^{\top}\textbf{v}_{i}\textbf{u}_{i}^{\top})^{\frac{1}{2}}\overset{(iii)}{=}\sigma_{i},$
@@
In this way, the inference pass is replaced by $\textbf{W}_{q}\textbf{X}+\textbf{L}_{A}\textbf{L}_{B}\textbf{X}$ for error reconstruction. Note that this technique is orthogonal to the weight-only quantization, which means ASER can be conducted based on the existing weight-only quantization methods, not limited to RTN. To sum up, the pseudo-code of our algorithm is as follows:
@@
13:     Compute S by Cholesky decomposition on $(\textbf{M}^{-1}\textbf{X})(\textbf{M}^{-1}\textbf{X})^{\top}$: $(\textbf{S}^{-1}\textbf{M}^{-1}\textbf{X})(\textbf{S}^{-1}\textbf{M}^{-1}\textbf{X})^{\top}=\textbf{I}$
14:     SVD decomposition: $\textbf{U}\mathbf{\Sigma}\textbf{V}^{\top}=\textbf{E}_{q}^{l}\textbf{S}$
@@
Experimental Details To ensure a fair comparison in our assessment, we standardize the usage of the calibration dataset for all quantization methods. All quantization is conducted by per-channel for weights and per-token for activations. We consistently employ a sample size of $128$ and maintain a uniform token sequence length of $2048$ for calibration throughout the evaluation process. All experiments are conducted on 4×4$\times$4 × NVIDIA A100 GPUs.
@@ Conclusion
In this work, from the perspective of model compression, we formulate and analyze the quantization error. Our study reveals that the error exhibits low-rank property, whose major components originate from the combined effect of outliers in both activations and weights. In response to these findings, we propose ASER, which utilizes whitening SVD to compensate for the integral error (Error Reconstruction), followed by analyzing and extracting activation outliers (Activation Smoothing). Experimental results demonstrate that ASER outperforms existing quantization algorithms on typical large language models in various quantization setups, with negligible theoretical overhead.